# IR Bounds on Theories with Spontaneously-Broken Lorentz Symmetry

### v2

@@ 3.2 Beyond the zero-gap approximation
Following the same steps as before one can prove that $N^{\mu}=Z^{\mu\nu}p_{\nu}$ still follows the geodesic equation Eq. (29), but now the geodesics are $S$-timelike,
@@
For our purposes this is the only change with respect to the previous case where the gap was negligible. Therefore the geometrical interpretation remains the same and the phase velocity can be defined as before by Eq. (32). Importantly, as a consequence of the presence of a gap we have $N^{\mu}p_{\mu}\neq 0$, meaning that in this case the speed defined through $N^{\mu}$ is no longer the phase velocity.
As a short and rather trivial example, we can consider a Lorentz invariant acoustic metric, i.e. $Z^{\mu\nu}=\eta^{\mu\nu}$ and $S_{\mu\nu}=\eta_{\mu\nu}$, in the presence of a gap $C^{2}$ (invariant mass $m^{2}$ in this case). The previous two equations give in this case, the dispersion relation
@@
Notice that with this, the quantity $\omega^{2}({\bf p})/{\bf p}^{2}$ is divergent for ${\bf p}\to 0$. Instead, the equation
$(N^{0})^{2}=m^{2}+{\bf N}^{2}\,,$ (45)
allows to derive the definition (32) for the ray velocity, which is well-defined at zero momentum. Given the form of the acoustic metric, the ray velocity takes the form:
@@
which for this specific case coincides with the group velocity. Equivalently, we find
$N^{0}=\frac{m}{\sqrt{1-{\bf v}^{2}}}\,,$ (47)
that is, ${\bf v}$ plays the role of velocity for a relativistic particle.
4 Bounds from the IR ray velocity
4.1 Conformal superfluid in D=3
We now return our focus to the low-energy ($\omega\ll\mu$) mode $\omega_{-}^{2}({\bf p})$, whose ray velocity will be momentum-dependent on account of the $c_{2,3}$ contributions. By bounding this velocity we can put constraints on these higher dimension coefficients. As discussed in the previous section, this being a gapped mode the proper velocity definition is through the ratio of the components of the ray vector $N^{\mu}$, as in Eq. (32).
From the dispersion relation (22) we can read off the acoustic metric [^6] Actually, we can only read it up to a conformal factor, which for our discussion is irrelevant so we choose $Z^{00}=1$.,
@@ 4.1 Conformal superfluid in D=3
with $Z^{\mu\nu}=diag(-1,\alpha^{2},\alpha^{2})$. An important remark here is that we are treating the contributions from HD operators, responsible for the ${\bf p}$-dependence in $\alpha^{2}({\bf p})$, as perturbative within the regime of validity of the EFT, implicitly giving a condition on $c_{2,3}$ and $p_{\mu}$. This prompts us define a $p_{\mu}$-dependent acoustic metric $Z^{\mu\nu}(p)$, which we then otherwise treat as usual. In particular, its inverse is simply $S_{\mu\nu}=diag(-1,1/\alpha^{2},1/\alpha^{2})$ and trading $p_{\mu}$ for $N^{\mu}=Z^{\mu\nu}(p)p_{\nu}$ can always be unambiguously done perturbatively. Rewriting the EoM in terms of $N^{\mu}$ we get
@@
which clearly would give the usual ${\bf v}^{2}=\alpha^{2}$ if there was no gap. Instead, in this case we have
@@
$\varepsilon\equiv\frac{{\bf p}\cdot{\bf N}}{m^{2}}=\alpha^{2}\frac{{\bf p}^{2}}{m^{2}}\,,$ (51)
@@
which leads to the following expression for the ray velocity
@@
As anticipated, the gap induced by the coupling between the scalar and gauge modes slows down the propagation. Notice also that this expression for the ray velocity reduces to the usual phase velocity for momenta well above the gap ($\varepsilon\gg 1$), i.e. ${\bf v}^{2}\simeq\omega_{-}^{2}({\bf p})/{\bf p}^{2}\simeq\alpha^{2}({\bf p})$. In particular, this becomes exact for the front velocity, the limiting velocity at infinite momentum which ultimately determines the support in spacetime of the retarded Green’s function, ${\bf v}^{2}(\infty)=\alpha^{2}(\infty)$.
We now proceed to impose our bound. As discussed, the expected behavior of the ray/phase velocity of a gapless mode in a conformal superfluid in $D=3$ is that it runs from the limiting value $c_{s}^{2}=1/2$ at zero momentum to $1$ at infinite momentum. Here however the gap has the effect of slowing down the mode to a crawl at momenta ${\bf p}^{2}\ll m^{2}$, instead allowing $\alpha^{2}({\bf p})$ in this regime to be practically unbounded. To provide a bound on $\alpha^{2}({\bf p})$, the physically meaningful condition is to demand the ray velocity ${\bf v}^{2}$ to remain below $c_{s}=1/2$ for sufficiently low momenta.
Taking this approach we demand,
@@
or, equivalently in terms of $\alpha^{2}({\bf p})$,
@@
On the other hand, we should not extend this condition too far in momentum, as it is only expected that eventually ${\bf v}^{2}$ will increase towards $1$. So a key question is just how far is far enough. A conservative answer is: far enough to bring the corresponding bound on $\alpha^{2}({\bf p})$ down to $1$, which is in turn the limit that we should impose at very high momenta, when ray velocity is indeed just phase velocity ${\bf v}^{2}\simeq\alpha^{2}({\bf p})$ and should approach $1$. The momentum ${\bf p}_{*}$ at which our bound above (54) becomes $\alpha^{2}({\bf p}_{*})\leq 1$ corresponds to $\varepsilon_{*}=1$, that is $\alpha^{2}({\bf p}_{*}){\bf p}_{*}^{2}=m^{2}$. Assuming moreover that the momentum dependence of $\alpha^{2}({\bf p})$ is always increasing, then we also have $\alpha^{2}({\bf p})\geq 1/2$ for all ${\bf p}$. Therefore, the pivot scale ${\bf p}_{*}$ lies in somewhere in the range
$m^{2}\leq{\bf p}_{*}^{2}\leq 2m^{2}\,.$ (56)
After this point, we instead just consider $\alpha^{2}({\bf p})\leq 1$ for momenta ${\bf p}^{2}\geq{\bf p}_{*}^{2}$.
To summarize our bound, we recall the definition for $\alpha^{2}$ from Eq. (22) and using Eq. (55), this entails
@@
with $\varepsilon\in[0,1]$. The optimal choice of $\varepsilon$ actually depends on the values of $\tilde{c}_{2}$ and $\tilde{c}_{3}$. In this form our result can be readily compared Ref. Creminelli:2022onn . For this purpose it is useful to change variables by defining $\frac{(1-\varepsilon)}{(1+\varepsilon)}\equiv 1-\xi^{2}$. For the range $\varepsilon\in[0,1]$ that we consider, the new variable is also $\xi\in[0,1]$, while the bound (57) can be rewritten as
@@
This confirms that our IR approach reproduces correctly the bounds derived in Creminelli:2022onn , see Eq. (58) there.
As a concluding remark, we can consider the limiting case where $b+d\to\infty$, where the vector field decouples. Recalling that $\tilde{c}_{2,3}=c_{2,3}/(b+d)$ and taking this limit in Eq. (57) at fixed $\varepsilon$, gives as a result a trivialization of the bound. In other words, we recover the expected outcome that for the pure scalar theory we are unable to extract a meaningful bound, as discussed in Sec. 2.1. From our current vantage point, this can be simply understood as a consequence of the vanishing of the gap in this limit which closes the window where the bound can effectively be applied.
4.2 Connection with $\langle JJ\rangle$ correlators
In Ref. Creminelli:2022onn analyticity bounds were derived for the conformal superfluid model studied above, by studying a Green’s function associated with the correlator of conserved currents $J^{\mu}$, computed in the case where the gauge field remains non-dynamical. The bounds obtained in this case coincide with our result in Eq. (54), which instead we obtained by studying the low-energy theory as is, allowing the gauge field to be dynamical and therefore kinetically mix with the scalar.
@@ 4.2 Connection with $\langle JJ\rangle$ correlators
where we rewrote the linear-in-$A_{\mu}$ term in the quadratic Lagrangian as
@@
At this point it is also clear that, if the gauge field $A_{\mu}$ where to be promoted to a dynamical field, its effective kinetic term after integrating out the scalar would be the quantity in curly brackets in the right-hand-side of Eq. (4.2), which is also the last expression in Eq. (61). This shows that the $\langle JJ\rangle$ correlator contains the information of the kinetic mixing between the scalar and the gauge field. Indeed, this correlator can potentially shift the pole of the $A_{\mu}$ propagator, which determines the kinematics of the gauge field excitations. In Fourier space
@@
where $\hat{J}^{\mu}(\omega,{\bf p})$ is $-i$ times the Fourier transform of $\tilde{J}^{\nu}(x,x^{\prime})$, with components given explicitly in Eqs. (11) and (12). Upon explicit computation, using a gauge fixing, the above can be seen to give the same condition as Eq. (18). This happens because the eigenmode of the kinetic matrix (10) mixes with $A_{\mu}$ and changes its propagation. Therefore the solution is precisely the mode $\omega_{-}^{2}({\bf p})$ given in Eq. (22), whose ray velocity we studied and where our bounds came from.
4.3 Example of UV completion
In this Section we outline a specific example of a UV completion of the conformal superfluid in $D=3$ in the absence of the gauge field, consisting of a conformal complex scalar field, as also exhibited in Ref. Creminelli:2022onn . This provides an explicit realization of the momentum dependence of the ray/phase velocity of the gapless mode, which transitions from $1/2$ at zero momentum to $1$ at momenta much larger than the scale of symmetry breaking $\mu$.
The Lagrangian for this UV-complete model is given by Badel:2019khk ,
@@ 4.3 Example of UV completion
where $\phi=\rho\,e^{i\theta}$ and $\lambda>0$. Upon integrating out the modulus field $\rho$ at tree level, the EFT action from Eq. (1) is obtained with coefficients Creminelli:2022onn
@@
plus higher-dimensional operators. Coupling to an external gauge field $A_{\mu}$ can be done by introducing a gauge-covariant derivative $\partial_{\mu}\to\nabla_{\mu}$. Expanding around a state with chemical potential $\mu$, $\theta=\mu t+\frac{\pi_{c}}{\sqrt{2}\bar{\rho}}$ and $\rho=\bar{\rho}+\frac{\delta\rho_{c}}{\sqrt{2}}$, where $\bar{\rho}\simeq\sqrt{\mu}/(3\lambda)^{1/4}$, the quadratic lagrangian for the fluctuations $\pi_{c}$ and $\delta\rho_{c}$ has the following kinetic matrix in Fourier space
$\mathcal{K}_{IJ}(\omega,{\bf p})=\begin{pmatrix}\omega^{2}-{\bf p}^{2}&2i\mu\omega\\ -2i\mu\omega&\quad\omega^{2}-{\bf p}^{2}-4\mu^{2}\\ \end{pmatrix}\,.$ (68)
The corresponding eigenmodes $\omega({\bf p})$ are obtained from $\det(\mathcal{K})=0$,
@@
The minus branch gives rise to a gapless mode, the superfluid phonon, while the plus branch describes a masive mode with $M^{2}=8\mu^{2}$. Within the regime of validity of the EFT, only the former can be excited. It is instructive to examine its dispersion relation in both the regimes of low and high momentum $|{\bf p}|$ with respect to $\mu$,
@@
Since this is a gapless mode, the ray and phase velocities coincide and can be read directly from ${\bf v}^{2}=\omega_{\phi-}^{2}({\bf p})/{\bf p}^{2}$. The above exhibits how the ray/phase velocity ${\bf v}^{2}$ of this mode interpolates between the soundspeed of a conformal superfluid in $D=3$, $c_{s}^{2}=1/2$, at low energies when the Lorentz invariance and the conformal invariance are spontaneously broken by the finite chemical potential $\mu$, and the speed of light $c^{2}=1$ in the unbroken phase at high energies.
5 Discussion

### v3

@@ III.2 Gapped and dispersive systems
Following the same steps as for the nondispersive case, one can still define the acoustic vector $N^{\mu}=Z^{\mu\nu}p_{\nu}$ and see that the EoM sets this vector as $S$-timelike:
@@
Of course, the vector $N^{\mu}$ still follows the geodesic equation, Eq. (43), although this is now trivial, as we are restricting to a flat acoustic metric, $\partial_{\rho}Z^{\mu\nu}=0$. Again, the velocity of rays can be defined as before through the acoustic vector $N^{\mu}$ by Eq. (37). This velocity will still coincide with the group velocity, as long as the acoustic metric is momentum independent, i.e., $\partial Z^{\alpha\beta}/\partial p_{\mu}=0$. If instead the metric depends on the momentum, we will have that the two notions of velocity are different:
$U_{g}^{\mu}=N^{\mu}+\frac{1}{2}\frac{\partial Z^{\alpha\beta}(p)}{\partial p_{\mu}}p_{\alpha}p_{\beta}\,.$ (48)
As a consequence of the presence of a gap, we now have $N^{\mu}p_{\mu}=-C^{2}\neq 0$, implying that in this case the speed defined through $N^{\mu}$ is no longer the phase velocity, as expected for a dispersive case.
As a short and rather trivial example, we can consider a Lorentz invariant acoustic metric, i.e. $Z^{\mu\nu}=\eta^{\mu\nu}$ and $S_{\mu\nu}=\eta_{\mu\nu}$, in the presence of a gap $C^{2}$ (invariant mass $m^{2}$ in this case). The previous two equations give the dispersion relation:
@@
Notice that with this, the phase velocity ${\bf v}_{p}^{2}=\omega^{2}({\bf p})/{\bf p}^{2}$ is divergent for ${\bf p}\to 0$. Instead, the equation
$(N^{0})^{2}=m^{2}+{\bf N}^{2}$ (50)
allows one to compute the velocity $\bf v$} from the acoustic vector $N^{\mu}$, Eq. (37), which is well-defined at zero momentum. Given the form of the acoustic metric in this example, it takes the form
@@
which coincides in this case with the group velocity ${\bf v}_{g}^{2}$, as discussed above.
While in this section we have focused in the case relevant to our purposes, of a space-time independent acoustic metric, let us remark, that in the case of a space-time dependent acoustic metric that depends on $p$, a generalization of Riemannian geometry is required and the relation between geodesics and auto-parallel curves becomes more involved [44].
IV Bounds in terms of the IR kinematics
IV.1 Conformal superfluid in D=3
We now return our focus to the low-energy ($\omega\ll\mu$) mode $\omega_{-}^{2}({\bf p})$, whose acoustic metric will be momentum-dependent on account of the $c_{2,3}$ contributions.
From the dispersion relation (25), we can read off the acoustic metric [^5] Actually, we can only read it up to a conformal factor, which for our discussion is irrelevant so we choose $Z^{00}=1$.
@@ IV.1 Conformal superfluid in D=3
with $Z^{\mu\nu}=diag(-1,\alpha^{2},\alpha^{2})$. An important remark here is that we are treating the contributions from HD operators, responsible for the $\bf p$}-dependence in $\alpha^{2}({\bf p})$, as perturbative within the regime of validity of the EFT, implicitly giving a condition on $c_{2,3}$ and $p_{\mu}$. This prompts us to define a $p_{\mu}$-dependent acoustic metric $Z^{\mu\nu}(p)$, which we then otherwise treat as usual. In particular, its inverse is simply $S_{\mu\nu}=diag(-1,1/\alpha^{2},1/\alpha^{2})$ and trading $p_{\mu}$ for $N^{\mu}=Z^{\mu\nu}(p)p_{\nu}$ can always be unambiguously done perturbatively. Rewriting the on-shell condition in terms of $N^{\mu}$, we get
@@
which clearly would give the usual ${\bf v}^{2}={\bf v}_{p}^{2}=\alpha^{2}$ if there was no gap. Instead, in this case, we have
@@
$\varepsilon\equiv\frac{{\bf p}\cdot{\bf N}}{m^{2}}=\alpha^{2}\frac{{\bf p}^{2}}{m^{2}}\,.$ (55)
@@
which leads to the following expression for the velocity $\bf v$} associated to the acoustic vector $N^{\mu}$,
@@
As anticipated, the gap induced by the coupling between the scalar and gauge modes slows down $\bf v$} at low momenta, where it significantly differs from the phase-velocity ${\bf v}_{p}$. Nevertheless, the two velocities approximately coincide for momenta well above the gap [^6] Note that $\varepsilon\gg 1$ can be in the regime of validity of the EFT as long as $b+d\gg c_{1}$, as we have that the mass gap is much below the cutoff of the EFT: $m^{2}\ll\mu^{2}$. ($\varepsilon\gg 1$), i.e. ${\bf v}^{2}\simeq\omega_{-}^{2}({\bf p})/{\bf p}^{2}\simeq\alpha^{2}({\bf p})$. Moreover, note that the $p$-dependence of the acoustic metric implies that $\bf v$} also differs from the group velocity ${\bf v}_{g}$.
As discussed, we expect the velocity of a gapless mode in a conformal superfluid in $D=3$ to run from the limiting value $c_{s}^{2}=1/2$ at zero momentum to $1$ at infinite momentum. As usual, we will assume that the running of this speed is strictly positive, corresponding to a stable system. Therefore, to provide a bound on $\alpha^{2}({\bf p})$, the physically meaningful condition should be to demand that velocities remain below $c_{s}^{2}=1/2$ for sufficiently low momenta, and below $c=1$ for high momenta. With this in mind, a relevant question is which notion of velocity makes it possible to capture the analyticity bound (7) through a sharp, momentum-independent condition. In the following, we show that such a momentum-independent condition can be found for the velocity $\bf v$} defined from the acoustic vector $N^{\mu}$.
According to the previous considerations, we examine a momentum-independent condition on $\bf v$}, namely,
@@
This can be equivalently written in terms of $\alpha^{2}({\bf p})$,
@@
This condition should not be imposed for too high momenta, as it is expected that eventually ${\bf v}^{2}$ will increase toward $1$. Since the coefficients of the EFT are not bounded when the theory is gapless, we know that the maximum momentum should scale like $m$, rather than $\mu$. For instance, we can stop at $\varepsilon=1$, [^7] From Eq. (55), we can see that momentum $\bf p$$_{*}$ at which $\varepsilon=1$ satisfies $\alpha^{2}({\bf p}_{*}){\bf p}_{*}^{2}=m^{2}$. Since $\alpha^{2}({\bf p})$ should always be increasing, then we also have $\alpha^{2}({\bf p})\geq 1/2$ for all $\bf p$}. Therefore, the scale $\bf p$$_{*}$ lies in somewhere in the range $m^{2}\leq{\bf p}_{*}^{2}\leq 2m^{2}$. When $\alpha^{2}({\bf p})$ saturates the bound (58), i.e. $\alpha^{2}_{\text{max}}=1$ at $\epsilon=1$, we have ${\bf p}_{*}^{2}=m^{2}$. where the momentum starts being important with respect to the mass in the energy balance of Eq. (25). As a matter of fact, we find that for $\varepsilon>1$, it will be $\alpha^{2}<1$, which limits the speed of excitations to be strictly below that of a Lorentz invariant particle at the same mass.  [^8] For a free, massive Lorentz-invariant particle, ${\bf v}^{2}={\bf v}_{g}^{2}\geq 1/2$ for $p^{2}\geq m^{2}$.
As we will see, these natural considerations lead to reproducing the analyticity bounds of Eq. (7), derived in [27]. Recalling the definition for $\alpha^{2}$ from Eq. (25) and using Eq. (59), we may now explicitly write Eq. (58) as
@@
with $\varepsilon\in[0,1]$. The optimal choice of $\varepsilon$ actually depends on the values of $\tilde{c}_{2}$ and $\tilde{c}_{3}$. In this form our result can be readily compared to the known analyticity bound (7). For this purpose, it is useful to change variables by defining $\frac{(1-\varepsilon)}{(1+\varepsilon)}\equiv 1-\xi^{2}$. For the range $\varepsilon\in[0,1]$ that we consider, the new variable is also $\xi\in[0,1]$, while (60) can be rewritten as
@@
This confirms that the bounds derived in Ref. [27] [see Eq. (58) therein], are in fact a sharp statement about $\bf v$}, that is, about the acoustic vector $N^{\mu}$.
Notice that the same bound cannot be expressed in terms of a momentum-independent condition on either the group or the phase velocity. In fact, as discussed in the previous Section, we know that in this situation $\bf v$} is different from both the phase velocity ${\bf v}_{p}$ and the group velocity ${\bf v}_{g}$. Therefore, the bound cannot constrain either of these velocities with a momentum-independent condition. This can be confirmed by direct computation,
${\bf v}_{p}^{2}$ $=$ $\frac{m^{2}}{{\bf p}^{2}}+\left[\frac{1}{2}+\bar{c}_{2}\left(-1+\frac{{\bf p}^{2}}{2m^{2}}\right)+\bar{c}_{3}\left(1+\frac{{\bf p}^{2}}{2m^{2}}\right)\right]\,$, (62)
${\bf v}_{g}$ $=$ $\left[\frac{1}{2}+\bar{c}_{2}\left(-1+\frac{{\bf p}^{2}}{m^{2}}\right)+\bar{c}_{3}\left(1+\frac{{\bf p}^{2}}{m^{2}}\right)\right]\frac{{\bf p}}{\omega}\,.$ (63)
It is also interesting to consider the limiting case where $b+d\to\infty$, where the gauge field decouples. Recalling that $\tilde{c}_{2,3}=c_{2,3}/(b+d)$ and taking this limit in Eq. (60) at fixed $\varepsilon$, gives as a result a trivialization of the bound. In other words, we recover the expected outcome that for the pure scalar theory we are unable to extract a meaningful bound, as discussed in Sec. II.1. From the point of view of our argument, this can be understood as a consequence of the vanishing gap for the scalar field, which implies that Eq. (58) can only be applied at zero momentum, giving a trivial bound.
IV.2 Interpretation
We have seen that the analyticity bounds can be transparently phrased as a momentum-independent condition (58) on the velocity $\bf v$} corresponding to the acoustic vector $N^{\mu}$, but due to dispersivity and the presence of a gap, this does not translate into an equally sharp bound on neither the phase velocity ${\bf v}_{p}$ nor the group velocity ${\bf v}_{g}$. It is useful then to see this condition directly as a constraint on the momentum of the gapped mode. Indeed, one may rewrite the bound of Eq. (58) as
$N^{\mu}\bar{p}_{\mu}\leq 0\,$, (64)
where $\bar{p}_{\mu}=\left(c_{s}|{\bf p}|,{\bf p}\right)$ is a fiducial on-shell momentum of a gapless particle with speed $c_{s}^{2}=1/2$, and $\bf p$} is the momentum of the gapped excitation (at which $N^{\mu}$ is evaluated). This means that the acoustic vector $N^{\mu}$ lies inside of the sound-cone of the gapless excitation. This soundcone can be defined through $\bar{N}^{\mu}=\bar{Z}^{\mu\nu}\bar{p}_{\nu}$, with $\bar{Z}^{\mu\nu}=diag(-1,c_{s}^{2},c_{s}^{2})$, as this vector satisfies: $\bar{N}^{\mu}\bar{p}_{\mu}=0$.
As discussed, the same condition cannot be phrased in a momentum independent way in terms of either the phase velocity $U_{p}^{\mu}$, the group velocity $U_{g}^{\mu}$, or correspondingly the momentum co-vector $p_{\mu}$, as the bound Eq. (58) would read $\bar{p}_{\mu}Z^{\mu\nu}(p)p_{\nu}\leq 0$, which is a condition on momentum dependent combinations of the components of $p_{\mu}$.
IV.3 Connection with $\langle JJ\rangle$ correlators
In Ref. [27] analyticity bounds were derived for the conformal superfluid model studied above, by studying a Green’s function associated with the correlator of conserved currents $J^{\mu}$, computed in the case where the gauge field remains non-dynamical. The bounds obtained in this case coincide with our result in Eq. (58), which instead we obtained by studying the low-energy theory as-is, allowing the gauge field to be dynamical and therefore kinetically mix with the scalar.
@@ IV.3 Connection with $\langle JJ\rangle$ correlators
where we rewrite the linear-in-$A_{\mu}$ term in the quadratic Lagrangian as
@@
At this point it is also clear that, if the gauge field $A_{\mu}$ were to be promoted to a dynamical field, its effective kinetic term after integrating out the scalar would be the quantity in curly brackets in the right-hand-side of Eq. (IV.3), which is also the last expression in Eq. (67). This shows that the $\langle JJ\rangle$ correlator contains the information of the kinetic mixing between the scalar and the gauge field. Indeed, this correlator can potentially shift the pole of the $A_{\mu}$ propagator, which determines the kinematics of the gauge field excitations. In Fourier space
@@
where $\hat{J}^{\mu}(\omega,{\bf p})$ is i-i times the Fourier transform of $\tilde{J}^{\nu}(x,x^{\prime})$, with components given explicitly in Eqs. (15) and (16). Upon explicit computation, using a gauge fixing, the above can be seen to give the same condition as Eq. (21). This happens because the eigenmode of the kinetic matrix (14) mixes with $A_{\mu}$ and changes its propagation. Therefore the solution to Eq. (71) is precisely the mode $\omega_{-}^{2}({\bf p})$ given in Eq. (25), which was central to the derivation of the bounds derived in Ref. [27] from the $\langle JJ\rangle$ correlator.
IV.4 Example of UV completion
In this Section we outline a specific example of a UV completion of the conformal superfluid in $D=3$ in the absence of the gauge field, consisting of a conformal complex scalar field, as also studied in Ref. [27]. This provides an explicit example of the momentum dependence of the velocity $\bf{v}$ obtained from the acoustic vector $N^{\mu}$ of the gapless mode, which transitions from $1/2$ at zero momentum to $1$ at momenta much larger than the scale of symmetry breaking $\mu$.
The Lagrangian for this UV-complete model is given by [45]
@@ IV.4 Example of UV completion
where $\phi=\rho\,e^{i\theta}$ and $\lambda>0$. Upon integrating out the modulus field $\rho$ at tree level, the EFT action from Eq. (1) is obtained with coefficients [27]
@@
plus higher-dimensional operators. Coupling to an external gauge field $A_{\mu}$ can be done by introducing a gauge-covariant derivative $\partial_{\mu}\to\mathcal{D}_{\mu}$. Expanding around a state with chemical potential $\mu$, $\theta=\mu t+\frac{\pi_{c}}{\sqrt{2}\bar{\rho}}$ and $\rho=\bar{\rho}+\frac{\delta\rho_{c}}{\sqrt{2}}$, where $\bar{\rho}\simeq\sqrt{\mu}/(3\lambda)^{1/4}$, the quadratic Lagrangian for the fluctuations $\pi_{c}$ and $\delta\rho_{c}$ has the following kinetic matrix in Fourier space
$\mathcal{K}_{IJ}(\omega,{\bf p})=\begin{pmatrix}\omega^{2}-{\bf p}^{2}&2i\mu\omega\\ -2i\mu\omega&\quad\omega^{2}-{\bf p}^{2}-4\mu^{2}\\ \end{pmatrix}\,$, (74)
where $I,J=\{\rho,\theta\}$. The corresponding eigenmodes $\omega({\bf p})$ are obtained from $\det(\mathcal{K})=0$,
@@
The minus branch gives rise to a gapless mode, the superfluid phonon, while the plus branch describes a massive mode with $M^{2}=8\mu^{2}$. Within the regime of validity of the EFT, only the former can be excited. It is instructive to examine its dispersion relation in both the regimes of low and high momentum $|{\bf p}|$ with respect to $\mu$,
@@
Since this mode is gapless, the phase velocity and the velocity from the acoustic vector $N^{\mu}$ coincide ${\bf v}^{2}={\bf v}_{p}^{2}=\omega_{\phi-}^{2}({\bf p})/{\bf p}^{2}$. The above example exhibits how the velocity ${\bf v}^{2}$ of this mode interpolates between the soundspeed of a conformal superfluid in $D=3$, $c_{s}^{2}=1/2$, at low energies when the Lorentz invariance and the conformal invariance are spontaneously broken by the finite chemical potential $\mu$, and the speed of light $c^{2}=1$ in the unbroken phase at high energies.
V Discussion